# Manipulating the Online Marketplace of Ideas

### v4

@@ Discussion
Social media platforms have enhanced the so-called attention economy, in which abundant content must compete for our scarce attention (45). But how to ensure that accurate, relevant, timely information wins this competition? To date, the policies that govern social media have been mainly guided by the concept of a free marketplace of ideas, rooted in John Milton’s centuries-old reasoning that truth prevails in a free and open encounter of opinions (46). In an ideal world, the wisdom of the crowd (47) would realize this vision by combining the opinions of many users (48). Unfortunately, several aspects of modern social media challenge the illusion of a public square or marketplace in which wise crowds access and select quality information (49, 50). First, information production is affected by information consumption (51), creating incentives to produce appealing but not necessarily high-quality content. Second, social influence undermines the wisdom of crowds because the information and opinions to which we are exposed online may be inauthentic, correlated through coordination (52, 53), or dominated by few influential individuals (54). Third, confirmation bias (55) can increase vulnerability to misinformation in social media (56, 11). Finally, the structure of information flow networks can distort perceptions and increase vulnerability to malicious actors (57, 58).
@@
A wealth of previous models and experiments have focused on information diffusion and popularity. Several studies have investigated the role played by network mechanisms affecting the popularity of individual posts, including exogenous and endogenous bursts of attention (59, 60), memory (61), novelty (62, 63), and position bias (64, 65). This literature considers the popularity of pieces of information in isolation. Market-like environments in which many messages compete for limited attention have received less consideration. Exceptions have considered the cost of learning about quality (66), distortions of quality assessments that result from aggregate knowledge of peer choices (52), and confirmation bias in the spread of misinformation (67). The SimSoM model proposed here extends the model of Weng et al. (68), who demonstrated that some posts inevitably achieve viral popularity irrespective of quality in the presence of competition among networked agents with limited attention. The model was formalized as a critical branching process and studied analytically, predicting that the popularity of posts follows a power-law distribution with heavy tails (69, 70, 71).
The current SimSoM model has several limitations. First, it assumes that all authentic agents are similar, when in reality users have diverse behaviors. For example, in the U.S., conservative social media users are more likely to share low-credibility content (7, 72, 44). Large language models have recently been proposed as a way to model agents with more realistic and heterogeneous behaviors (73). Second, SimSoM models information diffusion through simple contagion. While this approach is supported by empirical evidence (41, 74), complex contagion (25) can also play a role in the spread of certain types of harmful information (26). Finally, SimSoM neglects many mechanisms of actual online social media that may contribute to message exposure, such as search and the secret algorithmic details of each platform. Yet the model captures several universal ingredients of socio-technical networks, including selection criteria (appeal, social engagement, and recency) and limited individual attention. The empirical follower network also accounts for structural features (hubs, clustering, polarized communities) that play a key role in information diffusion (75, 76, 72, 77, 78). The model’s predictions are consistent with empirical findings about the difference in virality or lack thereof between low- and high-quality information on Twitter (9, 8). This suggests a reasonable balance between model realism and generality.
@@
Our results suggest that, surprisingly, inauthentic accounts do not need to target hubs (influentials); they can do more damage by connecting to random accounts. Future research should focus on whether the strategic placement of bad actors within/across polarized online communities can sway users toward a particular outcome, for example by distorting popularity perceptions (57, 58). Further work is also needed to characterize the effects of strategic targeting when the attacker has limited resources.
@@ Methods
SimSoM [^1] Code and data to implement the model and reproduce results are available at github.com/osome-iu/SimSoM is a parsimonious agent-based model inspired by the long tradition of representing the spread of ideas as an epidemic process where messages are passed along the edges of a network (90). The model simulates a directed follower network, as in Twitter/X, Mastodon, or Threads. Nodes represent agents (users) and links represent follower relations, which may or may not be reciprocal. The direction of a link goes from the follower to the followed (friend) account, capturing the flow of attention; when a friend’s post is reshared by a follower, information spreads in the opposite direction. In line with previous work (74, 41, 91), the diffusion process is modeled as simple contagion, where each exposure to a message results in the same resharing probability. We assume that the structure of the network is static (no unfollowing/blocking of accounts) during the information-spreading process. In contrast to classical epidemiological models, new messages are continuously introduced into the system in an exogenous fashion.
@@
We assume authentic agents prefer to reshare messages posted by their friends that are appealing, recent, and popular. This is based on empirical evidence that users are more likely to share popular content according to engagement signals (93). The probability that an agent shares a message from their news feed, allowing it to spread, is proportional to the message’s appeal, social engagement, and recency. More explicitly, let $M_{i}$ be the feed of $i$ ($|M_{i}|=\sigma$). The probability of message $m\in M_{i}$ being selected is $P(m)=a_{m}e_{m}r_{m}/\sum_{j\in M_{i}}a_{j}e_{j}r_{j}$ where $a_{m}$ is the appeal of message $m$, $e_{m}$ is the social engagement, i.e., the number of times it was (re)shared by $i$’s friends, and $r_{m}$ is the recency. Message recency decays with time as a stretched exponential function $r_{m}(t)=e^{-0.4t^{0.4}}$, where $t$ is the “age,” or the time passed since $m$ was first introduced to the agent’s news feed. This decay function is based on empirical online news engagement data (63). To model flooding, content from bad actors has exposure that is $\theta$ times higher than authentic content.
@@
On the other hand, we assume that bad actors strictly generate low-quality messages ($q=0$). The potentially deceptive nature of this content is modeled by the deception parameter $\phi$ ($0\leq\phi\leq 1$), the probability that a bad actor message is irresistibly appealing. With probability $\phi$, we set $a=1$, and with probability $1-\phi$ we draw appeal from the same distribution as for authentic accounts, $P(a)=(1+\alpha)(1-a)^{\alpha}$. If $\phi=0$, bad actors and authentic accounts generate messages with appeal drawn from the same distribution. If $\phi>0$, bad actor messages are more likely to have high appeal; the larger $\phi$, the greater the potential virality of low-quality content by bad actors.
Empirical Follow Network
We run simulations on a follower network derived from empirical Twitter data. This network was constructed from a 10% random sample of public tweets between June 1–30, 2017 (94). The data includes users (excluding likely automated accounts) that shared at least ten links to news sources, at least one of which was to a source labeled as low-quality. Only news sources with known political valence were considered. This sampling procedure captures a community of accounts that are both active and vulnerable to misinformation. Based on the shared links, most accounts in the dataset have an associated partisanship score (from $-1$ for left-leaning to $+1$ for right-leaning) defined as the average political bias of the news sources they share; and a misinformation score defined as the fraction of posts linking to low-credibility sources. Both partisanship and misinformation scores, included in the original network dataset, were based on news source labels from third-party fact-checkers (72).
From the original dataset, we select nodes with both partisanship and misinformation attributes. We further reduce the size of the network to speed up our simulations. We apply $k$-core decomposition to select $N=10{,}006$ nodes forming the $k=94$ core. Finally, we remove a random sample of edges to decrease the density of this core while preserving the average in/out-degree (number of friends/followers) of the original network ($k=0$ core). This results in $E=1{,}809{,}798$ edges; each node has on average approximately 180 friends/followers (Fig. 3, “hubs+communities” network).
Bad Actor Subnetwork
@@
The model assumes that inauthentic accounts follow each other to amplify low-quality content. To capture the characteristic presence of hubs and clustering (directed triads), the inauthentic subnetwork is created using a directed variant of the random-walk growth model (95). Specifically, the network is initialized with four fully connected nodes. We then add one new node at a time, assuming that each has fixed out-degree $k_{out}=3$. Once a new node $i$ comes into the network, it links to (follows) a randomly selected target node (friend) $j$. Each of the remaining $k_{out}-1=2$ friends are selected as follows: with probability expressed by a parameter $p$, $i$ follows a random friend of $j$’s; with probability $1-p$, $i$ follows another randomly selected node. Following friends of a friend has the effect of generating closed, directed triads and approximates a preferential attachment process, giving rise to hub nodes with high in-degree. The parameter $p$ thus models both hubs and clustering. We use $p=0.5$.
In addition, the bad actor subnetwork is designed to manipulate information flow and collective attention by spreading certain messages. Therefore we assume that bad actors get random authentic accounts to follow them: we add a directed link from each authentic node to each bad actor node with probability $\gamma$. (In the next subsection we present alternatives to this random targeting tactic.) The parameter $\gamma$ models the degree of infiltration of the network by bad actors (Fig. 2). When $\gamma=0$, there is no infiltration and bad actors are isolated, therefore harmless; the opposite extreme $\gamma=1$ indicates complete infiltration such that bad actors dominate the network. Because we are not concerned with the quality of messages on bad actor news feeds, they do not follow or reshare content from authentic agents for the analyses in this paper.
Bad Actor Targeting Tactics
@@
The system’s quality at each time step $t$ is calculated with an exponential moving average $\bar{Q}_{t}=\rho\bar{Q}_{t-1}+(1-\rho)Q_{t}$. As the simulation takes place, some of the messages become obsolete quickly, while others live longer and infect a larger fraction of the network. The simulation ends once the system reaches a steady state in which the difference between the average quality in two consecutive time steps is smaller than a threshold, i.e., $|\bar{Q}_{t}-\bar{Q}_{t-1}|/\bar{Q}_{t-1}<\epsilon$. The average quality reported for all analyses is calculated at the end of the simulation.
@@
For each set of parameters, we run 10 simulations starting from random conditions; the reported average quality is the average across many runs.
In scenarios where there are bad actors in the system, the default bad actor parameters are $\gamma=0.01$, $\phi=0$, and $\theta=1$. Except for the simulations exploring inauthentic targeting tactics, authentic followers of the bad actors are selected at random.
The parameters $\rho=0.8$ and $\epsilon=0.0001$ were tested to ensure that the system’s quality stabilizes at the steady state.
We shuffle the empirical follower network in various ways to derive the scenarios reported in Fig. 3. The “hubs+communities” network is the original one. In the “hubs” network, we shuffle the original network while preserving the degree distribution and destroying any community clustering. In the “communities” network, the outgoing endpoint of each edge in the original network is rewired to another node within the same community; this preserves the community structure and destroys the hub structure. In the “random” shuffle, all edges are rewired at random with uniform probability, destroying all hub and clustering structure.
@@ Acknowledgments
To estimate the relationship between quality and engagement, we start from a dataset of tweets about COVID-19 (96). The dataset consists of tweets containing the hashtags #coronavirus or #covid19 collected in 2020, selected from a 10% random sample of public tweets (97). From this set of tweets, we analyze only posts that shared links to news sources from March 9–29, 2020. Linked sources were extracted from URLs in the tweet metadata. URLs shortened with 70 popular URL shortening services were also expanded to reveal the actual sources. Posts with links to Twitter and other social media platforms were excluded. The remaining links were matched against sources with ratings obtained from NewsGuard333newsguardtech.com in April 2021. The final dataset contains 110,224 original Twitter posts that share at least one link with a NewsGuard rating. For each post we have both the rating of the shared source, used as a proxy for post quality, and the number of retweets. The latter is extracted from the metadata of the latest retweet of each post.
@@
Figure S4: Appeal distribution. (a) Distribution of appeal for authentic agent messages, for different values of the parameter $\alpha$. (b) Reshare count distribution of empirical posts, and of messages in the baseline simulation with only authentic accounts. The message reshare counts in the model are measured at the steady state.

### v5

@@ Discussion
Social media platforms have enhanced the so-called attention economy, in which abundant content must compete for our scarce attention (49). But how to ensure that accurate, relevant, timely information wins this competition? To date, the policies that govern social media have been mainly guided by the concept of a free marketplace of ideas, rooted in John Milton’s centuries-old reasoning that truth prevails in a free and open encounter of opinions (50). In an ideal world, the wisdom of the crowd (51) would realize this vision by combining the opinions of many users (52). Unfortunately, several aspects of modern social media challenge the illusion of a public square or marketplace in which wise crowds access and select quality information (53, 54). First, information production is affected by information consumption (55), creating incentives to produce appealing but not necessarily high-quality content. Second, social influence undermines the wisdom of crowds because the information and opinions to which we are exposed online may be inauthentic, correlated through coordination (56, 57), or dominated by few influential individuals (58). Third, confirmation bias (59) can increase vulnerability to misinformation in social media (60, 11). Finally, the structure of information flow networks can distort perceptions and increase vulnerability to malicious actors (61, 62).
@@
A wealth of previous models and experiments have focused on information diffusion and popularity. Several studies have investigated the role played by network mechanisms affecting the popularity of individual posts, including exogenous and endogenous bursts of attention (63, 64), memory (65), novelty (66, 67), and position bias (68, 69). This literature considers the popularity of pieces of information in isolation. Market-like environments in which many messages compete for limited attention have received less consideration. Exceptions have considered the cost of learning about quality (70), distortions of quality assessments that result from aggregate knowledge of peer choices (56), and confirmation bias in the spread of misinformation (71). The SimSoM model proposed here extends the model of Weng et al. (72), who demonstrated that some posts inevitably achieve viral popularity irrespective of quality in the presence of competition among networked agents with limited attention. The model was formalized as a critical branching process and studied analytically, predicting that the popularity of posts follows a power-law distribution with heavy tails (73, 74, 75).
The current SimSoM model has several limitations. First, it neglects many mechanisms of actual online social media that may contribute to message exposure, such as search, personalization, and other details of secret platform algorithms. For example, increasingly popular feed ranking algorithms are based less on what is shared by social connections and more on out-of-network recommendations. Bad-actor tactics could be affected differently by such algorithmic changes (see Supplementary Material). While the present model captures several universal criteria of feed algorithms (appeal, social engagement, recency), future work should investigate the vulnerability of different algorithmic affordances. Second, real user behaviors are driven by complex cognitive processes. Beyond finite attention, such cognitive aspects of information sharing are not explored here. Instead, the model assumes that all authentic agents follow a simple probabilistic rule in selecting the content to be shared. Large language models have recently been proposed as a way to model agents with more realistic behaviors (76). In addition, SimSoM models information diffusion through simple contagion. While this approach is in line with prior modeling (77), theoretical (45), and empirical (78) studies, complex contagion (25) can also play a role in the spread of certain types of harmful information (26). Despite these limitations, the current model’s predictions are consistent with empirical findings about the difference in virality between low- and high-quality information on Twitter (9, 8). This suggests a reasonable balance between model realism and generality.
@@
While the structure of online social networks evolves over time, the follower network used in our experiments provides a more realistic setting than synthetic networks to run our simulations. In particular, it captures features of real social networks —hubs, clustering, and political homophily structure— that have been observed consistently over the years and across platforms (83, 84, 85) and that are known to play key roles in information diffusion (86, 87, 88). Furthermore, our sampling procedure captures a community of accounts that are both active and vulnerable to misinformation, as needed to study manipulation tactics that target such communities. Finally, the assumption that the follower network is static during an information-spreading simulation is consistent with observations that the dynamics of unfollowing are slow (89).
Our results suggest that, surprisingly, inauthentic accounts do not need to target hubs (influentials); they can do more damage by connecting to random accounts. Future research should focus on whether the strategic placement of bad actors within/across polarized online communities can sway users toward a particular outcome, for example by distorting popularity perceptions (61, 62). Further work is also needed to characterize the effects of strategic targeting when the attacker has limited resources.
@@ Methods
SimSoM [^1] Code and data to implement the model and reproduce results are available at github.com/osome-iu/SimSoM is a parsimonious agent-based model inspired by the long tradition of representing the spread of ideas as an epidemic process where messages are passed along the edges of a network (97). The model simulates a directed follower network, as in Twitter/X, Mastodon, or Threads. Nodes represent agents (users) and links represent follower relations, which may or may not be reciprocal. The direction of a link goes from the follower to the followed (friend) account, capturing the flow of attention; when a friend’s post is reshared by a follower, information spreads in the opposite direction. In line with previous work (78, 45, 77), the diffusion process is modeled as simple contagion, where each exposure to a message results in the same resharing probability. We assume that the structure of the network is static (no unfollowing/blocking of accounts) during the information-spreading process. In contrast to classical epidemiological models, new messages are continuously introduced into the system in an exogenous fashion.
@@
We assume authentic agents prefer to reshare messages posted by their friends that are appealing, recent, and popular. This is based on empirical evidence that users are more likely to share popular content according to engagement signals (99). The probability that an agent shares a message from their news feed, allowing it to spread, is proportional to the message’s appeal, social engagement, and recency. More explicitly, let $M_{i}$ be the feed of $i$ ($|M_{i}|=\sigma$). The probability of message $m\in M_{i}$ being selected is $P(m)=a_{m}e_{m}r_{m}/\sum_{j\in M_{i}}a_{j}e_{j}r_{j}$ where $a_{m}$ is the appeal of message $m$, $e_{m}$ is the social engagement, i.e., the number of times it was (re)shared by $i$’s friends, and $r_{m}$ is the recency. Message recency decays with time as a stretched exponential function $r_{m}(t)=e^{-0.4t^{0.4}}$, where $t$ is the “age,” or the time passed since $m$ was first introduced to the agent’s news feed. This decay function is based on empirical online news engagement data (67). To model flooding, bad actors can be more active than authentic agents, sharing a message $\theta$ times such that it appears to have higher engagement than authentic content.
@@
On the other hand, we assume that bad actors strictly generate low-quality messages ($q=0$). The potentially deceptive nature of this content is modeled by the deception parameter $\phi$ ($0\leq\phi\leq 1$), the probability that a bad-actor message is irresistibly appealing. With probability $\phi$, we set $a=1$, and with probability $1-\phi$ we draw appeal from the same distribution as for authentic accounts, $P(a)=(1+\alpha)(1-a)^{\alpha}$. If $\phi=0$, bad actors and authentic accounts generate messages with appeal drawn from the same distribution. If $\phi>0$, bad-actor messages are more likely to have high appeal; the larger $\phi$, the greater the potential virality of low-quality content by bad actors.
Follower Network
We run simulations on a follower network derived from empirical Twitter data. This network was constructed from a 10% random sample of public tweets between June 1–30, 2017 (100). The data includes users (excluding likely automated accounts) that shared at least ten links to news sources, at least one of which was to a source labeled as low-quality. Only news sources with known political valence were considered. Based on the shared links, most accounts in the dataset have an associated partisanship score (from $-1$ for left-leaning to $+1$ for right-leaning) defined as the average political bias of the news sources they share; and a misinformation score defined as the fraction of posts linking to low-credibility sources. Both partisanship and misinformation scores, included in the original network dataset, were based on news source labels from third-party fact-checkers (84).
From the original dataset, we select nodes with both partisanship and misinformation attributes. We further reduce the size of the network to speed up our simulations. We apply $k$-core decomposition to select $N=10{,}006$ nodes forming the $k=94$ core. Finally, we remove a random sample of edges to decrease the density of this core while preserving the average in/out-degree (number of friends/followers) of the original network ($k=0$ core). This results in $E=1{,}809{,}798$ edges; each node has on average approximately 180 friends/followers (Fig. 3, “hubs+clusters” network).
Bad-Actor Subnetwork
@@
The inauthentic account subnetwork is created using a directed variant of the random-walk growth model (101). This captures the presence of hubs and clustering (directed triads). Specifically, the network is initialized with four fully connected nodes. We then add one new node at a time, assuming that each has fixed out-degree $k_{out}=3$. Once a new node $i$ comes into the network, it links to (follows) a randomly selected target node (friend) $j$. Each of the remaining $k_{out}-1=2$ friends are selected as follows: with probability expressed by a parameter $p$, $i$ follows a random friend of $j$’s; with probability $1-p$, $i$ follows another randomly selected node. Following friends of a friend has the effect of generating closed, directed triads and approximates a preferential attachment process, giving rise to hub nodes with high in-degree. The parameter $p$ thus models both hubs and clustering. We use $p=0.5$.
In addition, the bad-actor subnetwork is designed to manipulate information flow and collective attention by spreading certain messages. Therefore we assume that bad actors get random authentic accounts to follow them: we add a directed link from each authentic node to each bad-actor node with probability $\gamma$. (In the next subsection we present alternatives to this random targeting tactic.) The parameter $\gamma$ models the degree of infiltration of the network by bad actors (Fig. 2). When $\gamma=0$, there is no infiltration and bad actors are isolated, therefore harmless; the opposite extreme $\gamma=1$ indicates complete infiltration such that bad actors dominate the network. Because we are not concerned with the quality of messages on bad-actor news feeds, they do not follow or reshare content from authentic agents for the analyses in this paper.
Bad-Actor Targeting Tactics
@@
The system’s quality at each time step $t$ is calculated with an exponential moving average $\bar{Q}_{t}=\rho\bar{Q}_{t-1}+(1-\rho)Q_{t}$. As the simulation takes place, some of the messages become obsolete quickly, while others live longer and infect a larger fraction of the network. The simulation ends once the system reaches a steady state in which the difference between two consecutive values of the quality moving average is smaller than a threshold, i.e., $|\bar{Q}_{t}-\bar{Q}_{t-1}|/\bar{Q}_{t-1}<\epsilon$. See Supplementary Material for further details about model convergence to a steady state. The average quality reported for all analyses is calculated at the end of the simulation.
Cascade Construction
@@
For each set of parameters, we run 10 simulations starting from random conditions; the reported average quality is the average across these runs.
We explore values of the bad-actor parameters to cover broad ranges of manipulation tactics. The infiltration parameter $\gamma$ is a probability, and we consider values in the unit interval on a logarithmic scale to focus on more realistic low-probability values. The deception parameter $\phi$ also represents a probability and we examine the full unit interval. Finally, the flooding parameter $\theta$ is a multiplicative factor on the volume of posts, and we explore the range 1–64 on a logarithmic scale to capture the high volumes observed in empirical analyses (41). In scenarios with bad actors in the system, the default parameters are $\gamma=0.01$, $\phi=0$, and $\theta=1$. Except for the simulations exploring inauthentic targeting tactics, authentic followers of the bad actors are selected at random.
The parameters $\rho=0.8$ and $\epsilon=0.0001$ were tested to ensure that the system’s quality stabilizes at the steady state (see Supplementary Material).
We shuffle the follower network in various ways to derive the scenarios reported in Fig. 3. The “hubs+cluters” network is the original one. In the “hubs” network, we shuffle the original network while preserving the degree distribution and destroying any community clustering. In the “clusters” network, the outgoing endpoint of each edge in the original network is rewired to another node within the same cluster; this preserves the cluster structure and destroys the hub structure. In the “random” shuffle, all edges are rewired at random with uniform probability, destroying all hub and clustering structure.
Data Availability Statement
Code and data to implement the model and reproduce results are available at github.com/osome-iu/SimSoM.
@@ Acknowledgments
To estimate the relationship between quality and engagement, we start from a dataset of tweets about COVID-19 (102). The dataset consists of tweets containing the hashtags #coronavirus or #covid19 collected in 2020, selected from a 10% random sample of public tweets (103). From this set of tweets, we analyze only posts that shared links to news sources from March 9–29, 2020. Linked sources were extracted from URLs in the tweet metadata. URLs shortened with 70 popular URL shortening services were also expanded to reveal the actual sources. Posts with links to Twitter and other social media platforms were excluded. The remaining links were matched against sources with ratings obtained from NewsGuard333newsguardtech.com in April 2021. The final dataset contains 110,224 original Twitter posts that share at least one link with a NewsGuard rating. For each post, we have both the rating of the shared source, used as a proxy for post quality, and the number of retweets. The latter is extracted from the metadata of the latest retweet of each post.
@@
Figure S4: Appeal distribution. (a) Distribution of appeal for authentic agent messages, for different values of the parameter $\alpha$. (b) Reshare count distribution of messages in the baseline simulation of the model with only authentic accounts, and of empirical posts. The message reshare counts in the model are measured at the steady state.
S5 Convergence to steady state
Fig. S5(a) illustrates how the system reaches a steady state: after a transient period, the overall quality $\bar{Q}_{t}$ stabilizes around a consistent average value. To ensure that the overall quality of information $\bar{Q}$ is stable when the convergence criterion is met, we tested many values of the parameters $\rho$ and $\epsilon$. The time it takes to converge varies with different simulation parameters, like $\gamma$, $\phi$, and the population size. Using $\rho=0.8$ and $\epsilon=0.0001$, we observe that when the criterion is met, $\bar{Q}$ is stable for various values of the other parameters. Thus, we use these values in all the experiments reported in the main text.
Figure S5: (a) Overall quality across time steps for two simulation runs with different infiltration values. (b) Average quality at the end of prolonged simulations vs. when our convergence criterion is met. Plots represent the median (center bar), 50% (box) and 95% (whiskers) confidence intervals, and outliers (diamonds). The actual values from the simulations are represented as gray dots.
Additionally, we ran 20 simulations for an extended duration of 900 time steps, well beyond the typical convergence point. Fig. S5(b) compares the average quality $Q$ when the convergence criterion is met with that at the end of the extended simulations. The average quality values are similarly distributed in the two conditions, indicating stability. Note that the fluctuations of the overall quality $\bar{Q}$ are much smaller.
S6 Alternative feed algorithm
The model analyzed in the manuscript assumes that users are exposed to content (inventory) only form the accounts they follow. While this was typical of many social media platforms until recent years, in the wake of TikTok’s success several platforms now employ recommendation-based algorithms that mix in inventory from the entire network. While a full investigation of alternative algorithms is outside the scope of the present manuscript, we carried out a limited analysis to explore the effects of this kind of variations on our main results.
Figure S6: Comparison between the effects of individual bad-actor tactics on message quality, relative to the scenario without bad actors, with different feed algorithms. The top row shows results with the original algorithm (in-network inventory); the bottom row shows results with the alternative algorithm (50% in-network, 50% out-network). From left to right we report on the effect of varying infiltration $\gamma$ ($\theta=1$, $\phi$=0); flooding $\theta$ ($\gamma=0.01$, $\phi=0$); and deception $\phi$ ($\gamma=0.01$, $\theta=1$). Shading represents 95% confidence intervals across runs.
We consider a scenario in which 50% of each agent’s feed inventory comes from in-network posts (by friend accounts) and 50% from out-network posts (by the rest of the accounts). We repeat the analyses in Fig. 4(a,b,c), exploring the effects of bad-actor tactics based on infiltration, flooding and deception. We employ a smaller network obtained by down-sampling the follower network used in the main text ($N=1{,}000$ nodes, $E=30{,}030$ edges, 30 friends/followers per node on average). The results, shown in Fig. S6, are qualitatively similar to those obtained using the original algorithm based on only in-network inventory. However, infiltration has less impact while flooding and deception are more harmful in the mixed-inventory scenario. These changes can be understood by the observations that bad actors also receive and spread content from authentic agents, making their infiltration less harmful, while high-volume, high-appeal low-quality content can now spread farther. This analysis suggests that the harm caused by different bad-actor tactics can be mitigated or amplified by algorithmic affordances.